Because of EDC's massive expansion plan, their fixed costs are expected to skyrocket in the next year due to hiring, long-term asset purchases, and so on. The plant manager expects fixed costs to increase to $800,000. They sell desks for $700, and the variable costs for each are $45. Based on this information, your manager asks you to determine the following amounts:

  1. the break-even point in units,
  2. the break-even point in sales dollars, and
  3. the number of units needed to be sold for a profit of $200,000.
Answer #1

a) Break even Unit = Fixed cost/Contribution margin per unit = 800000/(700-45) = 1221 Units

b) Break even Sales = 1221*700 = $854700

c) Desired unit = (800000+200000)/655 = 1527 Units